OpenStudy (anonymous):

What is the distance between point A(-4, -2) and point B(5, 4) to the nearest tenth? A. 10.8 units B. 3.9 units C. 6.1 units D. 2.2 units

OpenStudy (anonymous):

@lexie<3 could you help me?

OpenStudy (anonymous):

i have the picture on the question on my test, i just dont know how to ppost it here

OpenStudy (anonymous):

For a mac you can do a screenshot by pressing shift command 3 gtg bai

OpenStudy (anonymous):

i dont have mac

OpenStudy (anonymous):

hmm well im thinking its between a and c

OpenStudy (anonymous):

i honestly have NO clue what it could be lol

OpenStudy (anonymous):

I think its c but I cant guarantee if its 100% right im sorry I cant help more

OpenStudy (anonymous):

its okay, when i turn in the test ill put the correct answer here :)

OpenStudy (anonymous):

ok

OpenStudy (trojanpoem):

(5, 4) - (-4, -2) = (9 , 6) distance = sqrt ( (9)^2 + (6)^2) = 10.81665 = 10.8 units